Given the following reaction

8 H+ + 5 Cu+ + MnO4- =====> 4 H2O + Mn+2 + 5 Cu 2+

Determine the weight percent Cu of a 1.84 gram sample if it takes 19.6 ml of 0.111 Molar MnO4- to titrate it. Find the moles of Cu first, then grams.
Answer #1

The balanced equation is 8H+ + 5 Cut + MnO4 =====> 4 H2O + Mn+2 + 5 Cu2+ Number of moles of MnO4- = M*V = 0.111 M * 19.6 mL =